# Catch Me if You Can: Detecting Unauthorized Data Use in Deep Learning Models

### v2

@@ 8. Conclusion and Future Work
This work presents MembershipTracker, a practical data provenance tool that can support ordinary users to audit whether their specific data are used to train deep learning models without their permission. MembershipTracker consists of a lightweight data marking process that can mark the target data with small and targeted changes, and a high-power set-based membership inference process that can reliably trace the provenance of the target data. By merely marking a small fraction of samples (0.005$\sim$0.1of the dataset), MembershipTracker effectively empowers the users to reliably audit the usage of their data (average 0% FPR@100% TPR), and it is also scalable to the large-scale (ImageNet) training setting. If their data are found to be misused, the users can take legal action or request the “right to be forgotten” in accordance with privacy regulations such as GDPR (gdp, [n. d.]) (e.g., based on techniques on machine unlearning (Bourtoule et al., 2021)).
@@
Second, future work can explore the potential of extending MembershipTracker to other domains like generative models. Large generative models are often trained on massive amount of data collected from the Internet (Rombach et al., 2022; Ramesh et al., 2022), and are the frequent targets where unauthorized use of personal/copyrighted contents arises (Dixit, 2023; Vincent, 2023). future work can investigate whether the large generative models would also be prone to memorize the samples marked with MembershipTracker under the billion-scale training environment.
Appendix A Model Training Details
For CIFAR10 and CIFAR100, we train a WideResNet-28-4 model on 25,000 samples for 100 epochs, with a learning rate of 0.1 (decayed by 5 at epoch 60, 80, 90), a weight decay of 5e-4 and momentum of 0.9. For TinyImageNet and ArtBench, we fine-tune an ImageNet-pretrained ResNet-18 model on 25,000 samples. For CelebA, we first follow (cel, [n. d.]) to create a 307-class subset from the dataset, and then similarly fine-tune an ImageNet-pretrained model on 2,000 samples. For fine-tuning, we use a small fine-tuning rate of 0.01 with 30 epochs and a momentum of 0.9. In all datasets except CelebA (where we find that fine-tuning on the entire set yields better accuracy), we use 20% of the training data as the validation set. We also use common data augmentations such as random cropping and horizontal flip.
Appendix B Detailed Results for Section 5.1
Earlier in Section 5.1 we evaluate MembershipTracker’s performance under different models and training-set sizes. The detailed results for this are in Table 6 and Table 7. In both cases, MembershipTracker is able to achieve 0% FPR@100% TPR with $<1\%$ accuracy drop the model.
Table 6. Evaluation on different models (parenthesized numbers show accuracy difference compared with the original models).
Model FPR@ Model
100% TPR accuracy
WideResNet 0.00 66.34 (-0.41)
ResNet 0.00 64.49 (-0.69)
ResNext 0.00 63.43 (-0.47)
DenseNet 0.00 67.14 (-0.16)
GoogleNet 0.00 69.09 (-0.33)
SeNet 0.00 65.26 (-0.37)
Table 7. Evaluation on different training-set sizes.
Training-set FPR@ Model
size 100% TPR accuracy
25,000 0.00 66.34 (-0.41)
20,000 0.00 63.73 (-0.16)
15,000 0.00 58.99 (-0.83)
10,000 0.00 50.28 (-0.09)
5,000 0.00 35.74 (-0.10)
Appendix C Ablation Study
@@
MembershipTracker consists of a data marking and set-based MI process, and we conduct a detailed ablation study into these two in Appendix C.1 and C.2, respectively.
C.1. Ablation of the Two-step Data Marking
@@
We first evaluate the effectiveness of the two data-marking components (Appendix C.1.1). Meanwhile, we also study how MembershipTracker’s performance varies under different number of target samples contributed by the users.
@@
Next, we evaluate how different data-marking intensities may affect MembershipTracker’s performance (Appendix C.1.2). Finally, we explore alternative data marking methods in Appendix C.1.3.
@@
(1) From top to the bottom: both 1ga1g and 1gb1g are crucial to MembershipTracker’s high auditing performance.
(2) From left to the right: MembershipTracker’s auditing performance degrades as the number of samples contributed by the users reduces. However, even in the challenging case where each user marks only 5 samples, the full MembershipTracker still incurs only 0.64% FPR@100% TPR - this highlights the importance of data marking in MembershipTracker.
@@ C.1.1. Investigating the two-step data marking process
Meanwhile, the performance difference between different approaches in Table 8 shrinks as the number of target samples contributed by the users increases. Thus, there is a low FPR even when auditing the original target samples (without data marking), e.g., 0.34% FPR@100% TPR on the setting with 15 target samples per user. This is in fact due to the effectiveness of the proposed set-based MI process [^5] The ablation study on MembershipTracker with and without the set-based MI process is in Appendix C.2. However, this is not enough and using the two-step data marking still achieves better performance. The performance advantage of performing data marking is even more profound under the more challenging cases with limited user samples (third and fourth column in Table 8).
(2) Performance variation under different number of target samples by the users. Table 8 also shows MembershipTracker’s performance degrades as the number of target samples by the users reduces (this trend is also similar to our earlier evaluation of the ImageNet training experiment in Section 6). There are two reasons.
@@
Figure 13. Performance evaluation under different image blending ratios (left) and noise perturbation budgets (right) (5 target samples per user). Zoom in to view the samples marked with different intensities.
C.1.2. Data marking with different parameters
@@
C.1.3. Alternative data marking methods
@@
C.2. Ablation of the Set-based MI Verification
@@
Figure 15. Comparing the technique by Wenger et al. (under varying data marking intensities) and MembershipTracker in terms of their MI effectiveness for data auditing.
Increasing the marking intensity in their approach can improve the MI success, but it: (1) still has significantly lower MI success than MembershipTracker; and (2) already incurs considerably higher distortion to the images than MembershipTracker (visualization in Fig. 16).
@@
As shown in Fig. 15, compared with the baseline approach without any data marking, the method by Wenger et al. can indeed increase the MI success. For instance, the green line in Fig. 15 (marked using $m=0.6$, which is equivalent to the default marking intensity in (Wenger et al., 2024)) increases the AUC from 0.7927 to 0.9253. However, this still has very limited performance in the low FPR regime. Increasing the marking intensity in their approach can further improve the MI success, but it: 1) comes at the cost of higher image distortion; and 2) still has lower auditing performance than MembershipTracker.
For instance, under $m=0.5$ (meaning 50% of the original features are replaced with the outlier features), their approach still yields much lower MI success than MembershipTracker (the gray Vs. blue line in Fig. 16), and the samples marked by their approach already become highly conspicuous compared with the original samples (the second and fifth column in Fig. 16). Using higher marking intensity ($m=0.3$) still suffers from a similar trade-off.
Our earlier discussion in Section 4.3.2 explains that injecting outlier features alone is not enough, as the original features hinder the model’s memorization on the outlier features. To overcome this, we propose a novel approach to inject a small amount of perlin noise, which works similar to adversarial perturbation and can significantly enhance the model’s memorization on the marked samples (as evaluated earlier in Appendix C.1.1) while preserving high image quality (the fourth column in Fig. 16). Combined with the proposed set-based MI process (another key component in our work), MembershipTracker is able to achieve perfect MI success (the blue line in Fig. 15) and outperforms the method by Wenger et al.
@@
Appendix E Comparison with Tramer et al. (Tramèr et al., 2022)
@@
For completeness, we evaluate how effective their approach is, if the users are able to mislabel their data for provenance purpose. We follow the targeted label flipping attack in Tramer et al., where for each target instance $(x,y)$, we inject multiple mislabeled data points $D_{\text{adv}}=\{(x,y^{\prime}),...,(x,y^{\prime})\}$ for some label $y^{\prime}\neq y$ (Tramèr et al., 2022). We inject different number of poisoned samples and compare their approach with MembershipTracker.
Table 10 shows the results, where the method by Tramer et al. indeed increases the exposure against the targeted samples, and the FPR@100% TPR is reduced from 49.33$\sim$56.32to 4.92$\sim$10.80%. Fig. 17 shows the ROC curve under the setting of 15 samples per user (similar trend on other settings). Meanwhile, the second row in Table 10 also shows that injecting more poisoned samples can contribute to higher MI success, which is also in line with (Tramèr et al., 2022). Nevertheless, even in the case where each user contributes 25 samples, the method by Tramer et al. still incurs 4.92% FPR@100% TPR; while MembershipTracker has 0% FPR@100% TPR.
Table 10. Comparing the technique by Tramer et al. and MembershipTracker in terms of their MI effectiveness for tracing data provenance.
Their approach works by injecting mislabeled samples to increase the MI success, which can reduce the FPR, but only to a limited extent.
In comparison, MembershipTracker does not require mislabeling any data, and still achieves considerably lower FPR.
Approach FPR@100% TPR
5 samples 15 samples 25 samples
(0.02% data) (0.06% data) (0.1% data)
Original 49.33 54.79 56.32
Tramer et al. 10.80 6.96 4.92
MembershipTracker 0.64 0.00 0.00
Figure 17. Comparing the performance between the approach by Tramer et al. and MembershipTracker (under 15 samples per user), and MembershipTracker achieves much higher MI success than their approach.
Appendix F Evaluating Huang et al. (Huang et al., 2024) under Limited Data Setting

### v3

@@ 7. Conclusion and Future Work
This work presents MembershipTracker, a practical data provenance tool that can support ordinary users to audit whether their specific data are used to train deep learning models without their permission. MembershipTracker consists of a lightweight data marking process that can mark the target data with small and targeted changes, and a high-power set-based membership inference process that can reliably trace the provenance of the target data. By merely marking a small fraction of samples (0.005$\sim$0.1of the dataset), MembershipTracker effectively empowers the users to reliably audit the usage of their data (average 100% TPR@0% FPR), and it is also scalable to the large-scale (ImageNet) training setting. If their data are found to be misused, the users can take legal action or request the “right to be forgotten” in accordance with privacy regulations such as GDPR (gdp, [n. d.]) (e.g., based on techniques on machine unlearning (Bourtoule et al., 2021)).
@@
Second, future work can explore the potential of extending MembershipTracker to other domains like generative models. Large generative models are often trained on massive amount of data collected from the Internet (Rombach et al., 2022; Ramesh et al., 2022), and are the frequent targets where unauthorized use of personal/copyrighted contents arises (Dixit, 2023; Vincent, 2023). Thus, future work can investigate whether the large generative models would also be prone to memorize the samples marked with MembershipTracker under the billion-scale training environment.
Appendix A Quantitative Evaluation of Image Visual Quality
A key goal of our work is to preserve high visual quality. While we provided some visualization examples earlier in Fig. 5, in this section, we conduct experiments to quantitatively assess the visual quality of the images marked by MembershipTracker.
We use three commonly used metrics: (1) Mean Squared Error (MSE), (2) Structural Similarity Index Measure (SSIM) and (3) Learned Perceptual Image Patch Similarity (LPIPS) to assess visual quality. MSE measures the average squared difference between corresponding pixel values of two images. SSIM is commonly used to assess the similarity between two images, considering changes in structural information, luminance, and contrast (Wang et al., 2004). LPIPS computes the distance between feature representations of two images extracted from a pre-trained neural network (Zhang et al., 2018b). This approach captures more nuanced differences in texture and structure, aligning more closely with human perception of image quality.
We use the built-in SSIM implementation from TorchMetric (tor, [n. d.]). We adopt the LPIPS implementation from the original authors (Zhang et al., 2018b), and we follow their technique to use an AlexNet for extracting the feature representations. We use 5,000 random images for evaluation under three settings: 1g11g images marked with full MembershipTracker (with OOD feature and procedural noise); 1g21g images blended with OOD feature (without procedural noise); and 1g31g the strawman approach (which replaces the target images with OOD feature). We report the results in Table 8.
Table 8. Quantitative evaluation of image visual quality. 1g11g Images marked with full MembershipTracker (with OOD feature and procedural noise); 1g21g images blended with OOD feature (without procedural noise); and 1g31g the strawman approach (which replaces the target images with OOD feature).
Settings MSE SSIM LPIPS
1g11g 0.0077 0.8424 0.0244
1g21g 0.0072 0.8900 0.0173
1g31g 0.0814 0.0923 0.4076
For full MembershipTracker ( 1g11g), the average SSIM is high (0.8424), and the LPIPS and MSE are low (0.0244 and 0.0077), indicating high similarity between the marked and original images. Removing the procedural noise ( 1g21g) can improve the visual quality, under which the MSE is reduced to 0.0072, the SSIM is increased to 0.8900 and LPIPS reduced to 0.0173. Compared with 1g11g and 1g21g, the strawman approach ( 1g31g) has the highest MSE and LPIPS and the lowest SSIM, which is understandable as it replaces the target image with OOD features. Overall, our results quantitatively validate that the images marked by full MembershipTracker can still preserve high visual quality under different metrics.
Appendix B Ablation Study
@@
MembershipTracker consists of a data marking and set-based MI process, and we conduct a detailed ablation study into these two in Appendix B.1 and B.2, respectively. As explained earlier in Section 5.2, we use FPR@100% TPR as the evaluation metric.
B.1. Ablation of the Two-step Data Marking
@@
We first evaluate the effectiveness of the two data-marking components (Appendix B.1.1). Meanwhile, we also study how MembershipTracker’s performance varies under different number of target samples contributed by the users.
@@
Next, we evaluate how different data-marking intensities may affect MembershipTracker’s performance (Appendix B.1.2). Finally, we explore alternative data marking methods in Appendix B.1.3.
B.1.1. Investigating the two-step data marking process
@@
Meanwhile, the performance difference between different approaches in Table 9 shrinks as the number of target samples contributed by the users increases. Thus, there is a low FPR even when auditing the original target samples (without data marking), e.g., 0.34% FPR@100% TPR on the setting with 15 target samples per user. This is in fact due to the effectiveness of the proposed set-based MI process [^6] The ablation study on MembershipTracker with and without the set-based MI process is in Appendix B.2. However, this is not enough and using the two-step data marking still achieves better performance. The performance advantage of performing data marking is even more profound under the more challenging cases with limited user samples (third and fourth column in Table 9).
(2) Performance variation under different number of target samples by the users. Table 9 also shows MembershipTracker’s performance degrades as the number of target samples by the users reduces (this trend is also similar to our earlier evaluation of the ImageNet training experiment in Section 5.3). There are two reasons.
@@
Figure 13. Performance evaluation under different image blending ratios (left) and noise perturbation budgets (right) (5 target samples per user).
B.1.2. Data marking with different parameters
@@
B.1.3. Alternative data marking methods
@@
B.2. Ablation of the Set-based MI Verification
@@
Figure 15. Comparing the technique by Wenger et al. (under varying data marking intensities) and MembershipTracker in terms of their MI effectiveness for data auditing. MembershipTracker achieves much higher TPR than their approach in the low FPR regime.
@@
As shown in Fig. 15, compared with the baseline approach without any data marking, the method by Wenger et al. can indeed increase the MI success. For instance, the green line in Fig. 15 (marked using $m=0.6$, which is equivalent to the default marking intensity in (Wenger et al., 2024)) increases the AUC from 0.7927 to 0.9253. However, this still has very limited performance in the low FPR regime (e.g., 0% TPR@0.01% FPR). Increasing the marking intensity in their approach can further improve the MI success, but it: 1) comes at the cost of higher image distortion; and 2) still has lower auditing performance than MembershipTracker.
For instance, under $m=0.5$ (meaning 50% of the original features are replaced with the outlier features), their approach still yields much lower MI success than MembershipTracker, e.g., 3.8% TPR@0.01% FPR vs. 100% TPR by MembershipTracker (the gray Vs. blue line in Fig. 16), and the samples marked by their approach already become highly conspicuous compared with the original samples (the second and fifth column in Fig. 16). Using higher marking intensity ($m=0.3$) still suffers from a similar trade-off.
Our earlier discussion in Section 4.1.2 explains that injecting outlier features alone is not enough, as the original features hinder the model’s memorization on the outlier features. To overcome this, we propose a novel approach to inject a small amount of perlin noise, which works similar to adversarial perturbation and can significantly enhance the model’s memorization on the marked samples (as evaluated earlier in Appendix B.1.1) while preserving high image quality (the fourth column in Fig. 16). Combined with the proposed set-based MI process (another key component in our work), MembershipTracker is able to achieve perfect MI success (the blue line in Fig. 15) and outperforms the method by Wenger et al.
@@
Appendix D Comparison with Tramer et al. (Tramèr et al., 2022)
@@
For completeness, we evaluate how effective their approach is, if the users are able to mislabel their data for provenance purpose. We follow the targeted label flipping attack in Tramer et al., where for each target instance $(x,y)$, we inject multiple mislabeled data points $D_{\text{adv}}=\{(x,y^{\prime}),...,(x,y^{\prime})\}$ for some label $y^{\prime}\neq y$ (Tramèr et al., 2022). We consider 100 target users where each user injects 15 poisoned samples. We compare their approach with MembershipTracker next.
Figure 17. Comparing the performance between the approach by Tramer et al. and MembershipTracker. MembershipTracker achieves much higher TPR than their approach in the low FPR regime.
Fig. 17 compares the performance by both approaches. As shown, the method by Tramer et al. indeed increases the exposure against the targeted samples. For example, under a 0.1% FPR threshold, the TPR is increased from from 0% to 36%. Nevertheless, the increase of MI success for the method by Tramer et al. is still highly limited, while MembershipTracker can achieve 100% TPR@0.1% FPR.
Therefore, compared with the approach by Tramer et al., MembershipTracker (1) represents a much more practical data provenance tool that does not require mislabeling any data, and (2) still achieves significantly higher MI success for tracing data provenance.
Appendix E Evaluating Huang et al. (Huang et al., 2024) under Limited Data Setting